All angle are in degrees 
 
 
Cos20+cos40+cos60-4cos10.cos20.cos30 is equal to??????

Grade:11

1 Answers

Aditya Gupta
5 years ago
Cos20+cos40=2cos30cos10
4cos10.cos20=2cos30+2cos10
so, 4cos10.cos20cos30=2[cos30+cos10]cos30
so, subtracting
Cos20+cos40 – 4cos10.cos20cos30=2cos30[cos10 – (cos30+cos10)]= –2cos^230= –3/2
adding cos60 to both sides
Cos20+cos40 – 4cos10.cos20cos30+cos60= –3/2+1/2 =  – 1